# Sparse Multivariate Linear Regression with Strongly Associated Response Variables

### v2

@@ Abstract
We propose new methods for multivariate linear regression when the regression coefficient matrix is sparse and the error covariance matrix is dense. We assume that the error covariance matrix has equicorrelation across the response variables. Two procedures are proposed: one is based on constant marginal response variance (compound symmetry), and the other is based on general varying marginal response variance. Two approximate procedures are also developed for high dimensions. We propose an approximation to the Gaussian validation likelihood for tuning parameter selection. Extensive numerical experiments illustrate when our procedures outperform relevant competitors as well as their robustness to model misspecification.
@@ 1 Introduction
Similar approaches that jointly estimate the error precision matrix and the regression coefficient matrix have been proposed. For instance, Lee and Liu [6] assumed that the response variables and the predictors have a joint multivariate normal distribution, and Wang [14] decomposed the multivariate regression problem into a series of penalized conditional log-likelihood of each response conditional on the predictors and other responses. Navon and Rosset [7] extended this to multivariate linear mixed effects models, and Chang and Welsh [3] considered an extension of Rothman et al. [12]’s method to multivariate robust linear regression. Zhu [17] considered a convex reparametrization of Rothman et al. [12]’s joint optimization problem. These methods assumed that the error precision matrix is sparse, which may be unreasonable in some applications with highly correlated response components. In addition, in high-dimensional settings, the graphical lasso subproblem [4] used in the block-wise coordinate decent algorithm of Rothman et al. [12] struggles when the error precision matrix is dense. These dense cases call for values of the penalty tuning parameter near zero, which lead to algorithm failures or very long computing times when solving the graphical lasso subproblem.
This motivated us to develop a new multivariate linear regression method for high-dimensional settings that works when the error precision matrix is dense. Similarly to Rothman et al. [12], we jointly estimate the regression coefficient matrix and the error precision matrix by minimizing the negative Gaussian loglikelihood. However, unlike their approach, we assume the error covariance matrix has compound symmetry (or equicorrelation). Our primary goal is to estimate the regression coefficient matrix when the responses are highly correlated. Although our assumed error correlation structure is simple, our estimation of the regression coefficient matrix is robust to misspecification of the error correlation. We propose an efficient computational algorithm to compute our estimators and we also propose an approximation to the validation likelihood for tuning parameter selection.
@@ 2 Problem setup and proposed estimator
where $\Omega(\eta^{2},\theta)=\left[\eta^{2}\left\{\theta 1_{q}1_{q}^{\prime}+(1-\theta)I_{q}\right\}\right]^{-1}$ and $\lambda\in[0,\infty)$ is a tuning parameter. Similarly to Rothman et al. [12], the $L_{1}$ penalty on $B$ encourages sparse regression coefficient estimation. We label the solution to this optimization as the Multivariate Regression with Compound Symmetry [MRCS] estimator. When the response components are on different scales, we define an alternative estimator described in Section 4.
@@ 3.1 Algorithm for exact computation
By the derivation illustrated in A of our supplementary material, the penalized estimator in (3) is equivalent to the following:
@@ 3.2 Algorithm for approximate computation
When $p\geq n$, the residual sample covariance matrix, $(Y-XB)^{T}(Y-X)/n$, is singular. When there exists $\bar{B}$ such that any one of the columns of $Y-X\bar{B}$ is zero, then the objective function value for updating $\Omega$ can be made arbitrarily negative by increasing the corresponding diagonal element of $\Omega$ sufficiently, which implies that a global minimizer does not exist. In addition, the computational time of the blockwise coordinate descent for fitting both regression coefficients and the error precision matrix tends to be significantly higher in increased dimensions of $p,\,q$ [12]. To improve computational efficiency and ensure the well-posed optimization problem in high dimensions, we provide an approximate solution to our procedure by following the idea of approximate MRCE proposed in Rothman et al. [12]. As in Algorithm 1, we compute the initial $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Then, for each $\lambda$, we compute $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ by (8). After this step for (inverse) covariance update, we compute the proposed approximate solution $\hat{B}$ with known $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ by (9). We suppress the dependence of $\hat{B}$ and $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ on $\lambda$ for notational simplicity. We refer the approximate solution to (4) as ap.MRCS. We summarize the procedure in Algorithm 2 below.
@@ 4 Extension to general equicorrelation covariance matrix
and $\lambda\in[0,\infty)$ is a tuning parameter. Note that the problem (11) is bi-convex in $B$ and ($(\{\eta_{i}\}_{i=1}^{q},\theta)$. As in Algorithm 3.2, we treat ($(\{\eta_{i}\}_{i=1}^{q},\theta)$ as the first block and $B$ as the second block in blockwise coordinate descent.
For a fixed $\hat{B}^{(k)}$, we solve for ($(\{\eta_{i}\}_{i=1}^{q},\theta)$ cyclically. By simple algebra (see A in our supplementary material), we compute the update for $\eta_{j}$ by
@@
For efficiency, we do not require full convergence in each inner loop over ($(\{\eta_{i}\}_{i=1}^{q},\theta)$. Instead, we conduct single-iteration updates for each $\{\eta_{i}\}_{i=1}^{q}$ and $\theta$ by (12) and (13).
@@
  For each fixed value of $\lambda$, initialize $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Set $k=0$ and ($(\{\hat{\eta}_{i}^{(0)}\}_{i=1}^{q},\hat{\theta}^{(0)})=(1_{q},0)$.
  Step 1: Compute one-step update of ($(\{\hat{\eta}_{i}^{(k+1)}\}_{i=1}^{q},\hat{\theta}^{(k+1)})(\hat{B}^{(k)})$ by solving (12)–(13) cyclically.
@@ 4.1 Approximate solutions II
As in Section 3.2, we propose an approximate solution to (11). Again, we start the algorithm with $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$, the combined lasso initializer. Then, for each $\lambda$, we compute ($(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})$ from (11). However, the difference between this approximation and the canonical MRGCS is that we conduct the updating step for ($(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})$ until convergence instead of a single-iteration update. This subproblem is obviously convex in ($(\{\eta_{i}\}_{i=1}^{q},\theta)$. Next, we compute the final $\hat{B}$ by solving (9) once. We call this approximation ap.MRGCS. We summarize it in Algorithm 4 below.
@@
  Step 1: Compute ($(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})(\hat{B}^{(0)})$ cyclically by solving (12) and (13) until convergence.
  Step 2: Compute $\hat{B}(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})$ by solving (9).
@@ 5 Asymptotic statistical properties
We study large sample asymptotic behavior of our estimator MRCS defined by (4) using the same asymptotic framework as Lee and Liu [6]. We study high-dimensional statistical properties of ap.MRCS (see Algorithm 3.2) in C of the supplementary material. We keep $p$ and $q$ fixed with $n>p+q$ throughout this section. As stated in Proposition 1 of Zou [18], the equally-weighted lasso-penalized estimator does not possess the oracle property. Thus, to ensure the oracle property, it is encouraged to assign weights, $w_{jk}$, to each $|B_{jk}|$ in the penalty term. We let $w_{jk}=1/|B^{({\rm ols})}_{jk}|^{r}$ ($r>1$) be the weight for $|B_{jk}|$, where $B^{({\rm ols})}_{jk}$ is $(j,k)$-th element of the regression coefficient matrix obtained from ordinary least squares. Using these weights, the slightly modified optimization (than (4)) we consider in this section is as follows:
@@
    : There exists $(\tilde{\eta}^{2},\tilde{\theta})$ $\sqrt{n}$-consistent estimator of $(\eta_{*}^{2},\theta_{*})$.
@@
In the following theorem, we provide the oracle guarantee, the limit distribution of $\hat{B}$, and the joint limit distribution of $(\hat{\eta}^{2},\hat{\theta})$. This result is analogous to Theorem 3 of Lee and Liu [6].
@@
Under the conditions (A1)(A4), assuming that $\lambda\sqrt{n}\rightarrow 0$ and $\lambda n^{(r+1)/2}\rightarrow\infty$, then, there exists a local minimizer $(\hat{B},\hat{\eta}^{2},\hat{\theta})$ to the optimization problem (14) that satisfies
@@
where $W=(1,-1/\eta_{*}^{2})^{T}$, and $V$ is defined in B of the supplementary material (see (27)).
@@ 6 Tuning parameter selection
We note that we do not use $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ from the output of MRCS. Rather we use the one-step estimators used for ap.MRCS that are computed in the Step 1 of Algorithm 3.2. We use the same $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ for both MRCS and ap.MRCS. This is because the stability of ap.MRCS output turned out to be significantly better than that of the canonical MRCS. Furthermore, we use $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ which assumes compound symmetry error covariance even when we fit MRGCS or ap.MRGCS which assume varying marginal error variances. This is because the output of ap.MRCS was significantly more stable in the tuning parameter selection compared with that of ap.MRGCS as well as the MRGCS output. We provide further support regarding this tuning parameter selection procedure through extensive simulations in G.1 and G.2 of the supplementary material.
@@ 7 Simulations
where $[k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$; #$\#A$ stands for the number of elements in the set $A$.
@@ 7.1 Setting I: Constant $\eta$
In this section, we consider $(p,q)\in\{(20,50),(50,20),(80,80)\}$, We refer the readers to H.1 in the supplementary material for the results of low dimensional simulations ($p=q=20$). We vary $s_{1}\in\{0.1,0.5\},\,s_{2}\in\{0.1,0.5,1\},\theta\in\{0,0.5,0.75,0.9,0.95\}$, and fix $\eta_{i}=1$ for all $i=1,\ldots,q$. We drop MRCE due to its substantially high computation time. When $p\geq n$ we exclude MRCS, MRGCS from the set of competitors, where we still consider ap.MRCS, and ap.MRGCS, since they can avoid the residual covariance instability (see Section 3.2). The results for each setting are based on 50 independent replications. Throughout the results, the trend across the estimators in the prediction error is nearly equivalent to that in the model error. Hence, we mainly discuss the model error as our primary criterion.
@@ 7.1.1 Results when $(p,q)=(20,50)$
The complete simulation results are in Figure 1922. As $\theta$ increases, all our equicorrelation based estimators consistently outperforms ap.MRCE as well as two lasso methods. The best non-oracle estimators were MRCS and ap.MRCS. MRGCS and ap.MRGCS also showed competitive performance. In TNR, sep.lasso showed the best rate in general, MRCS and ap.MRCS outperformed MRGCS and ap.MRGCS, and ap.MRCE was the worst. On the other hand, in TPR, two lasso solutions were generally the worst, and all joint optimization methods showed increasing TPR with $\theta$. This indicates the joint optimization methods tend to provide more dense solution than lasso as $\theta$ gets increased.
@@ 7.1.2 Results when $(p,q)=(50,20)$
[…]

### v3

@@ Abstract
We propose new methods for multivariate linear regression when the regression coefficient matrix is sparse and the error covariance matrix is dense. We assume that the error covariance matrix has equicorrelation across the response variables. Two procedures are proposed: one is based on constant marginal response variance (compound symmetry), and the other is based on general varying marginal response variance. Two approximate procedures are also developed for high dimensions. We propose an approximation to the Gaussian validation likelihood for tuning parameter selection. Extensive numerical experiments illustrate when our procedures outperform relevant competitors as well as their robustness to a moderate degree of model misspecification.
@@ 1 Introduction
Similar approaches that jointly estimate the error precision matrix and the regression coefficient matrix have been proposed. For instance, Lee and Liu [7] assumed that the response variables and the predictors have a joint multivariate normal distribution, and Wang [12] decomposed the multivariate regression problem into a series of penalized conditional log-likelihood of each response conditional on the predictors and other responses. Navon and Rosset [8] extended this to multivariate linear mixed effects models, and Chang and Welsh [3] considered an extension of Rothman et al. [11]’s method to multivariate robust linear regression. Zhou et al. [14] and Chan et al. [2] assumed that the covariance structure of the response variables follows a factor model with latent factors. Zhu [15] considered a convex reparametrization of Rothman et al. [11]’s joint optimization problem. These methods assumed that the error precision matrix is sparse, which may be unreasonable in some applications with highly correlated response components. In addition, in high-dimensional settings, the graphical lasso subproblem [4] used in the block-wise coordinate decent algorithm of Rothman et al. [11] struggles when the error precision matrix is dense. These dense cases call for values of the penalty tuning parameter near zero, which lead to algorithm failures or very long computing times when solving the graphical lasso subproblem.
This motivated us to develop a new multivariate linear regression method for high-dimensional settings that works when the error precision matrix is dense. Similarly to Rothman et al. [11], we jointly estimate the regression coefficient matrix and the error precision matrix by minimizing the negative Gaussian loglikelihood. However, unlike their approach, we assume the error covariance matrix has compound symmetry (or equicorrelation). Our primary goal is to estimate the regression coefficient matrix when the responses are highly correlated. Although our assumed error correlation structure is simple, our estimation of the regression coefficient matrix is robust not only to a moderate degree of misspecification in the error correlation but also to non-sparse structures in the true regression coefficients when the error covariance is correctly specified. We propose an efficient computational algorithm to compute our estimators and we also propose an approximation to the validation likelihood for tuning parameter selection.
@@ 2 Problem setup and proposed estimator
where $\Omega(\eta^{2},\theta)=\left[\eta^{2}\left\{\theta 1_{q}1_{q}^{\prime}+(1-\theta)I_{q}\right\}\right]^{-1}$ and $\lambda\in[0,\infty)$ is a tuning parameter. Similarly to Rothman et al. [11], the $L_{1}$ penalty on $B$ encourages sparse regression coefficient estimation. We label the solution to this optimization as the Multivariate Regression with Compound Symmetry [MRCS] estimator. When the response components are on different scales, we define an alternative estimator described in Section 4. The MRCE method [11] replaces $\Omega$ in equation (3) with a general error precision matrix consisting of the $q(q+1)/2$ free parameters that are not parametrized by $\eta$ and $\theta$. Added to the $L_{1}$-penalty on the regression coefficient, it adopts an additional $L_{1}$-penalty on the off-diagonal entries of $\Omega$ to encourage sparsity in the precision matrix. Additional simulation results demonstrating the computational failure of MRCE under dense error covariance structures (e.g., compound symmetry) are provided in Section 2 of the Supplementary material.
@@
Remark 1.
An anonymous referee pointed out that our compound symmetry–aware estimator reduces to a special case of methods proposed by Zhou et al. [14] and Chan et al. [2], which assume that the covariance structure of the response variables follows a factor model, when the model has a single latent factor with a factor loading vector of ones. In Section 4 of the supplementary material, we show that, when the compound symmetry structure is correctly specified, our method outperforms the factor model with the oracle number of latent factors while requiring lower computational cost. Moreover, because the number of latent factors is typically unknown in practice, our computational advantage becomes increasingly significant as the number of latent factors increases. Thus, our model can be seen as a special case of factor model, which uses more parameters to model the error structure more flexibly. However, as we discuss in Section 4 of the supplementary material, this added complexity combined with the nonconvex nature of the optimization often cause numerical difficulties in practice.
@@ 3.1 Algorithm for exact computation
By the derivation illustrated in A, the penalized estimator in (3) is equivalent to the following:
@@
The reparameterization converts the original non-convex optimization problem in $(\eta,\theta)$ into a separable problem in $(\alpha,\gamma)$, where each subproblem becomes a strictly convex scalar optimization admitting closed-form updates. Thus it improves numerical stability and accelerates convergence while automatically obeying the positivity and ordering constraints. Just as importantly, this alternative ($\alpha,\gamma$) formulation also facilitates a more streamlined derivation of the large-sample limit distribution of ($\hat{\eta},\hat{\theta}$) in Theorem 1.
@@ 3.2 Algorithm for approximate computation
When $p\geq n$, since the residual sample covariance matrix, $(Y-XB)^{T}(Y-XB)/n$, is singular or near-singular, the canonical MRCS algorithm’s alternating updates can encounter difficulties. For example, when there exists $\bar{B}$ such that any one of the columns of $Y-X\bar{B}$ is zero, then the objective function value for updating $\Omega$ can be made arbitrarily negative by increasing the corresponding diagonal element of $\Omega$ sufficiently, which implies that a global minimizer does not exist. In addition, the computational time of the blockwise coordinate descent for fitting both regression coefficients and the error precision matrix tends to be significantly higher in increased dimensions of p,qp,\,q [11].
To improve computational efficiency and ensure the well-posed optimization problem in high dimensions, we provide an approximate solution to our procedure by following the idea of approximate MRCE [ap.MRCE] proposed in Rothman et al. [11]. ap.MRCE addresses these issues by first obtaining an initial sparse estimate of $\Omega$ and then performing a single update of $B$, with simulations demonstrating its competitive performance. Similarly, our approximate algorithm limits the number of updates to reduce instability and improve computational speed, while preserving accurate estimation of $B$.
As in Algorithm 1, we compute the initial $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Then, for each $\lambda$, we compute $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ by (8). After this step for (inverse) covariance update, we compute the proposed approximate solution $\hat{B}$ with known $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ by (9). We suppress the dependence of $\hat{B}$ and $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ on $\lambda$ for notational simplicity. We refer the approximate solution to (4) as ap.MRCS. We summarize the procedure in Algorithm 2 below.
@@ 4 Extension to general equicorrelation covariance matrix
and $\lambda\in[0,\infty)$ is a tuning parameter. As in Algorithm 3.2, we treat $(\{\eta_{i}\}_{i=1}^{q},\theta)$ as the first block and $B$ as the second block in blockwise coordinate descent.
For a fixed $\hat{B}^{(k)}$, we solve for $(\{\eta_{i}\}_{i=1}^{q},\theta)$ cyclically. By simple algebra (see A), we compute the update for $\eta_{j}$ by
@@
For efficiency, we do not require full convergence in each inner loop over $(\{\eta_{i}\}_{i=1}^{q},\theta)$. Instead, we conduct single-iteration updates for each $\{\eta_{i}\}_{i=1}^{q}$ and $\theta$ by (12) and (13).
@@
 For each fixed value of $\lambda$, initialize $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Set $k=0$ and $(\{\hat{\eta}_{i}^{(0)}\}_{i=1}^{q},\hat{\theta}^{(0)})=(1_{q},0)$.
Step 1: Compute one-step update of $(\{\hat{\eta}_{i}^{(k+1)}\}_{i=1}^{q},\hat{\theta}^{(k+1)})(\hat{B}^{(k)})$ by solving (12)–(13) cyclically.
@@ 4.1 Approximate solutions II
[…]
We conducted additional simulations comparing MRCS, MRGCS, and their approximate versions under two initializers, combined and separate lasso, using a much finer grid of tuning parameters (see Section 1.3 of the supplementary material). Results show that initialization has little impact on MRCS and ap.MRCS, whereas MRGCS and ap.MRGCS are sensitive under high sparsity, where separate lasso performs better. Based on the favorable performance in our simulation settings and the computational simplicity, we adopted the combined lasso initializer. However, in practice, we recommend practitioners to evaluate both initialization strategies and choose the one that achieves the best performance via cross-validation.
@@
Step 1: Compute $(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})(\hat{B}^{(0)})$ cyclically by solving (12) and (13) until convergence.
[…]
@@ 5 Asymptotic statistical properties
We study large sample asymptotic behavior of our estimator MRCS defined by (4) using the same asymptotic framework as Lee and Liu [7]. We keep $p$ and $q$ fixed with $n>p+q$ throughout this section. As stated in Proposition 1 of Zou [16], the equally-weighted lasso-penalized estimator does not possess the oracle property. Thus, to ensure the oracle property, it is encouraged to assign weights, $w_{jk}$, to each $|B_{jk}|$ in the penalty term. We let $w_{jk}=1/|B^{({\rm ols})}_{jk}|^{r}$ ($r>1$) be the weight for $|B_{jk}|$, where $B^{({\rm ols})}_{jk}$ is $(j,k)$-th element of the regression coefficient matrix obtained from ordinary least squares. Using these weights, the slightly modified optimization (than (4)) we consider in this section is as follows:
@@
[…]
@@
[…]
@@
[…]
@@
where $W=(1,-1/\eta_{*}^{2})^{T}$, and $V$ is defined in B (see (27)).
@@ 6 Tuning parameter selection
We note that we do not use $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ from the output of MRCS. Rather we use the one-step estimators used for ap.MRCS that are computed in the Step 1 of Algorithm 3.2. We use the same $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ for both MRCS and ap.MRCS. This is because the stability of ap.MRCS output turned out to be significantly better than that of the canonical MRCS. Furthermore, we use $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ which assumes compound symmetry error covariance even when we fit MRGCS or ap.MRGCS which assume varying marginal error variances. This is because the output of ap.MRCS was significantly more stable in the tuning parameter selection compared with that of ap.MRGCS as well as the MRGCS output. We provide further support regarding this tuning parameter selection procedure through extensive simulations in Section 1.1 and 1.2 of the supplementary material [5].
@@ 7 Simulations
where $k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$; #A\#A stands for the number of elements in the set $A$.
@@ 7.1 Setting I: Constant $\eta$
In this section, we consider $(p,q)\in\{(20,50),(50,20),(80,80)\}$, We refer the readers to Section 3.1 in the supplementary material [5] for the results of low dimensional simulations ($p=q=20$). We vary $s_{1}\in\{0.1,0.5\},\,s_{2}\in\{0.1,0.5,1\},\theta\in\{0,0.5,0.75,0.9,0.95\}$, and fix $\eta_{i}=1$ for all $i=1,\ldots,q$. We drop MRCE due to its substantially high computation time. When $p\geq n$ we exclude MRCS, MRGCS from the set of competitors, where we still consider ap.MRCS, and ap.MRGCS, since they can avoid the residual covariance instability (see Section 3.2). The results for each setting are based on 50 independent replications. Throughout the results, the trend across the estimators in the prediction error is nearly equivalent to that in the model error. Hence, we mainly discuss the model error as our primary criterion.
@@ 7.1.1 Results when $(p,q)=(20,50)$
Complete simulation results are provided in Figure 26–29 of the supplementary material [5]. As $\theta$ increases, all equicorrelation-based estimators consistently outperform ap.MRCE and the two lasso methods. Among non-oracle estimators, MRCS and ap.MRCS performed best, with MRGCS and ap.MRGCS also competitive. For TNR, sep.lasso achieved the highest rates, followed by MRCS and ap.MRCS, while ap.MRCE performed worst. In contrast, for TPR, the lasso methods were generally poorest, and all joint optimization methods achieved higher TPR as $\theta$ increased, indicating their tendency to provide denser solutions.
@@ 7.1.2 Results when $(p,q)=(50,20)$
Complete simulation results are in Figure 30–33 of the supplementary material [5]. ap.MRCS was the best non-oracle estimator, with ap.MRGCS performing similarly except for $(s_{1},s_{2})=(0.5,0.1)$, where ap.MRCE outperformed it when $\theta=0$ and $0.95$. Overall, our non-oracle estimators outperformed ap.MRCE and the lasso methods. The TNR/TPR patterns were similar to those for $(p,q)=(20,50)$, though ap.MRCE showed improved TNR.